OpenStudy (anonymous):

Algebra 2 help please!!! I really am in need of help, will FAN and Medal!!!

OpenStudy (anonymous):

Create an equation. Use the graph below to create the equation of the rainbow parabola.

OpenStudy (anonymous):

@theopenstudyowl

OpenStudy (anonymous):

Someone please help

OpenStudy (anonymous):

I know I should use the equation y=a(x-h)^2+k, but what would I put into the equation?

OpenStudy (thebadyboy):

hello

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

hmmm.... im not sure have you ever dona a question like this before

OpenStudy (thebadyboy):

i hate algebra

OpenStudy (anonymous):

I've done a question like this before, it's just a confusing topic.

OpenStudy (anonymous):

All I know so far is that I have to create an equation using the graph and I can start by using y=a(x-h)^2+k and substituing the vertex (0, 36) for h and k.

OpenStudy (anonymous):

@MAli13chineta

OpenStudy (anonymous):

I think I figured most of it out :In order to create an equation using the graph, I can start by using y=a(x-h)^2+k and substituting the vertex (0, 36) for h and k and a point (6, 0) for x and y. This will create the equation 0=a(6-0)^2+36. Now, I just have to solve for a: 0=a(6-0)^2+36 >>> 0=a(6)^2+36 >>> 0=a(36)+36 >>> 0=a

OpenStudy (anonymous):

Well actually for this I wouldn't have to solve for a, so the equation would just be 0=a(6-0)^2+36 right?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hello! :)

OpenStudy (misty1212):

vertex is ((0,36)\) and it opens down so it must be \[y=ax^2+36\] for some negative number \(a\)

OpenStudy (misty1212):

since it is zero at \(-6\) and \(6\) you can pretty much guess that it is \[y=-x^2+36\]because that has the two zeros you want

OpenStudy (anonymous):

Wait, I'm a bit confused. How did you get y=ax^2+36?

OpenStudy (misty1212):

or you could multiply out \[a(x-6)(a+6)=a(x^2-36)\] then replace \(x\) by \(0\), set the result equal 36 and solve \[a(0-36)=36\] for \(a\) and see that \(a=-1\)

OpenStudy (misty1212):

ok lets go slow

OpenStudy (misty1212):

you see from the picture that the vertex is \((0,36)\) right?

OpenStudy (anonymous):

For this problem, I don't have to solve for a, I just need to create an equation and then use it for the next 2 questions about the graph

OpenStudy (anonymous):

Because after I create an equation, I have to answer this: Create a table of values for a linear function. A drone is in the distance, flying upward in a straight line. It intersects the rainbow at two points. Choose the points where your drone intersects the parabola and create a table of at least four values for the function. Remember to include the two points of intersection in your table.

OpenStudy (misty1212):

i just looked at your solution above, and it is right you got \[y=-x^2+36\]which is correct

OpenStudy (anonymous):

Oh ok, thank you!! can you help me with the second question pertaining to the graph which is: Create a table of values for a linear function. A drone is in the distance, flying upward in a straight line. It intersects the rainbow at two points. Choose the points where your drone intersects the parabola and create a table of at least four values for the function. Remember to include the two points of intersection in your table.

OpenStudy (anonymous):

Here's the graph:

OpenStudy (anonymous):

@misty1212 ??